Suppose when real disposable income is​ $5,000, planned real consumption is​ $4000. When real disposable income increases to​ $6000, planned real saving increases by​ $500. The new planned real consumption expenditures is

A. ​$5,000. B. ​$4,500. C. ​$6,000. D. ​$3,500.
Answer #1

Real savings =real disposable income -real consumption

=5000-4000=$1000

After increase real disposable incomei to $6000

Real saving increases by=$1000+500=$1500

New real consumption =new real disposable income-real savings

=$6000-$1500=$4500

Therefore the answer is option B.$4500